FEL Upgrade Project Weekly Brief
February 3-7, 2003
 
 
Highlights:
This week installation and qualification activities continued as we are nearing completion of the first light configuration for the FEL Upgrade project. The last magnet and vacuum chamber was installed in the first recirculation arc and the wiring hook-up was completed on the second optical cavity assembly (the output coupler).

Commissioning Activities:
Beam transport:
The second GX magnet and its vacuum chamber were installed in the first recirculation arc yesterday. This completes the electron beam transport hardware in this recirculation arc. The last major vacuum element in the first light configuration is the 4th vacuum chamber in the 4 magnet (GW) optical chicane assembly. This chamber is undergoing final vacuum qualification and should be released shortly.

David Douglas and his graduate student, Chris Tenant, qualified the bends in the first arc. The magnets meet specification within his criteria.

Optics:
The wiring was completed for the second optical cavity assembly (the output coupler) and check-out of this wiring has started. The first light optic (for 6 microns) and a cavity length measurement optic (800nm) are installed in the High Reflector optical cavity assembly.

Photoelectron Gun:
The gun assembly was disassembled to look for problems which occurred during last week's high voltage conditioning to 240 kV. A vacuum leak was found in the main large conflat flange and was repaired. The ceramic stack is leak tight. We will reassemble the ceramic stack for a full voltage test with the gun power supply while we analyze and fix arcing damage on the electrode coating (see details below).

Management:
F. Dylla and W. Skinner met with ONR Project Management (G. Graff and J. Albertine) on Feb. 6 to review the cost schedule and technical progress on the FEL Upgrade and to discuss the proposed statement of work for the remainder of FY03.

F. Dylla attended the quarterly meeting of the Virginia Research and Technology Advisory

Commission on Feb.4-5 that included a presentation to the Virginia Congressional delegation on upcoming R&D priorities.

WBS 4 (Injector):
The electrode structure was disassembled and inspected to allow a determination of what caused the low voltage field emission in the gun assembly last week. We found damage to the field emission coating near the cathode that may have been caused by field emission from the cracked cathode earlier. The damage is concentrated close to the cathode position and not the high field point of the cathode ball. A few arc spots were also found at the high field point of the cathode support tube and the coating on the support tube was bubbling and delaminated in spots. Silica dust from the exfoliated coating was found near the cathode ball. The group of people who coated the electrodes has been assembled to study the damage and determine the cause of the coating damage. We are working on setting up to recoat the electrodes using a recipe that previously demonstrated the best results. A test run with a test tube will be done first to qualify the recipe. In parallel with work on the electrode, the leak in the stack was isolated to a flange seal and repaired. The stack will now be assembled without the electrode and run up the full voltage of the high voltage power supply to qualify the ceramic stack and the high voltage power supply. We will also be testing previously coated electrodes to eliminate some other possible causes of coating failure.

WBS 8 (Instrumentation):
Many hours were spent this week attempting to attach the 4 and 8 micron aluminum foils to the arc paddle beam viewers. Our efforts have been made extremely difficult by the tragic passing of our colleague Karel Capek last summer. Since then the division has been lacking a diagnostics technician. Both skill of a craftsman and the touch of an artist are required to fabricate and install the diagnostics that are so critical to the successful commissioning a new accelerator.

The layout and installation has begun for the Danfysik (super trim - 20 Amp) power supplies, Apogee interfaces and control modules. These will power the sextupole lights and the QT's. Software addressing was assigned for the each of these magnets. Local testing with multiple Apogee VME modules has shown no interrupt conflicts. DC power for the 1st Arc String has been completed. DC power was also completed for the optical klystron. Commissioning shake out of the Injector, Linac trims and quads was completed this week (thanks EES). G. Biallas and D. Douglas will be providing test plans for completing the commissioning of the 1st Arc String, Injector, 2nd Arc String power supply ( temporary dump) and DY.

Drawings were completed for the Fabrication and Assembly of the HVPS Interlock Chassis and the Dump Water Temperature/Conductivity Monitoring Chassis. The drawings for the Dump Water Pressure/Temperature Interface Chassis are being revised to make them easier to read and reflect changes made to the chassis.

A new Dump Current Buffer Driver PCB is being developed. The schematic is complete and boards will be ordered next week. This is based on the DRV134 digital audio differential line driver. This had been tested last year and proved very successful. These will be used on both the fixed and insertable dumps.

A new 4 channel closed loop stepper motor control chassis was brought on line this week. Our standard 8 channel stepper motor driver chassis was modified for the Oregon Micro Systems VME58-4E (E for encoder). The chassis has 4 stepper drivers and 4 encoder connections. These are for the optical cavity mirror change motors. In the past we have assumed no slippage of requested steps but that is not good enough for positioning the optical cavity mirrors.

Video System: In the testing out the new 64:16 crosspoint chassis, a component failure surfaced that was only apparent for the I/O address that the box was set to receive at. After, this was discovered, we were quickly able to bring the 3rd switcher chassis into full operation. The video system has gone from 128 channel to 192 working channels. The channels are segmented as follows: 1....32: Beam viewers, 33...64: Multi-slits and ALL other electron beam video including UV viewers, 65...96: User Lab Video Inputs (6 Labs x 5 Channels per Lab), 97...112: All Area Cameras, 113..128: Zone Patches (Zones 1..5, Clean room Control Room & OCR) 2 channels per, 129..152: Optical Transport Video & including diagnostic patches, 153..160: FEL Vault Half-Rack Patch Panels, 161..192: 32 channels left for future expansion. Also, extensive clean-up and re-cabling has been done to the video (FL02B09) rack. Pictures of FL02B09 all cleaned-up is this weeks feature at http://laser.jlab.org/wbs8

The Engineering Department Manager has signed the FEL ODH Risk Assessment. This places the vault and gallery areas in an ODH 0 condition during normal operations and whenever the cryomodule is connected to the helium header. Any other conditions will need to be addressed separately by a TSOP as appropriate. Security maintenance was necessary for our PCs (laser.jlab.org and felvideopc.jlab.org) that are accessible off-site.

WBS 9 (Transport):
Work Arounds
Substitution of IR Demo 180° Bend (DY) for the Upgrade 180° Bend (GY) in the First Arc.
The vacuum chamber was attached to the new Chambers and fit well.
Dipoles
Optical Chicane Dipoles (GW)
No progress.
Injector Dipoles (DU/DV)
No action
Arc 180 Degree Dipoles (GY)
Don Bullard was not available to work on the set-up of this magnet because of his work on the gun. As an alternate, the Jlab Shop assigned the machinist who has been helping the magnet test area. However, he had some critical tasks to be done for the first light configuration and will not be available until early next week. In the interim, the shop is fixing the interference of the field clamp over the coil crossover wings on the left side as well as making field clamp adjuster brackets of a newer design that worked out so well on the GX/GQ series. Jim will fix the problem of the GG coil leads, where the conductor was cut off short, by counter boring the bulkhead style bracket. We also discovered that those fittings were wrongly specified as steel, which is not compatible with our low conductivity water. Replacement fittings are on order.
Wang is working on potting the second coil of the second GY. They predict shipment in two weeks.
Arc Bend & Reverse Bend Dipoles (GQ, GX)
Very good news, David Douglas and his graduate student, Chris Tenant, qualified the bends in the first arc after doing field integrals along the nominal trajectories and checks for field flatness in a series of transverse cuts the size to the required beam path. The magnets meet specification within his criteria.
We reached agreement with Wang NMR to make a replacement GX coil for the #2 GX. Magnet. They will air freight it.
Wang NMR shipped the remaining two GQ magnets on Wednesday. They should arrive on next Thursday.
The magnet measurement plan changed due to pressing Physics work. The group will take up measurement of the remaining 5 Arc dipoles in two weeks.
Quadrupoles
Trim Quad (QT)
The modified trim coils to create the GC vertical corrector function within the four trim quads arrived and were inserted into two QT quads.
Sextupole (SF)
No progress
Octupole (OT)
The bids are in evaluation.
Beam Line and Vacuum
Our shop finished the last chamber in Optical Chicane and the vacuum group baked it to remove some oil contamination. It will be in by Monday morning.
The magnets of the 1st Arc were wired to the Arc 1 power supply.
The two 180* Bend chambers continue to await the receipt of the diagnostic windows from Ceramaseal. The latest projected delivery date is late February.
The Vacuum Group connected chambers in the first Arc.
Remaining correctors and quadrupoles the Linac were tested for powering with their power supplies.

WBS 11 (Optics):
On the HR, the water cooling loop of the HR was connected, two mirrors (the 3" deformable mirror assembly (DMA) and 2" 800 nm) installed, and wiring extensions added. This afternoon we'll finish water connections and add the 220 L/s ion pump. The laser safety shutter was installed, as well as the smaller 40 L/s ion pump. We'll start a dry nitrogen purge that will run over the weekend, then early next week pump to a range where we can start the ion pumps. On the OC, all internal wiring is complete and will be checked out early next week.

One insertable mirror was built, installed on it's cross, and tested with a vacuum load in both the vertical and horizontal positions. No binding was noted and the repeatability appears excellent. We will install it as soon as we complete the OC.

Other commissioning activities:
The new EO modulator replacement (a chopper with a single small hole) was successfully tested. As configured, it can produce 400 us long pulses at 2 Hz with nearly 100% transmission. This can be used to replace one of the EO modulators if we wish more laser throughput, as well as a somewhat simpler EO optimization routine. We still need to determine the best way to control it with EPICS. We used the stretched output of our ultrafast laser system in a pump-probe arrangement to test the performance our electro-optic sampling setup. We assisted the gun team by loaning them our optical inspection microscope for support tube inspection.